HVAMS03: Logbook Entries

HVAMS03: Site 5 Messages, 11 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
137 Mon 20-Oct-2003st5: Goes Swapped5militzer
121 Wed 15-Oct-2003St 5: routine5delany
117 Mon 13-Oct-2003St5: kh2oV checked5delany
105 Tue 07-Oct-2003Site 5 visit, Oct 75holden
81 Mon 29-Sep-2003Site 5 visit, Sept 295oncley
67 Thu 25-Sep-2003Site 5 visit, Sept 245horst
54 Mon 22-Sep-2003Site 5 visit, Sept 205horst
47 Thu 18-Sep-2003Site 5 visit; Sept 175horst
31 Fri 12-Sep-2003Visit: kh2o cleaned...5militzer
29 Fri 12-Sep-2003Site Visit: CARD problem, Kh20 installed, HFT fixed5militzer
19 Mon 11-Aug-2003Site 5 CR10X calibration file5semmer


137: GOES, Site 5, Mon 20-Oct-2003 08:56:59 EDT, st5: Goes Swapped
10/19/03 13:35-14:55

Station was 'down' per GOES data.
Visited site: still running but old SE110 transmitter at 2W output.
Swapped in repaired SE120 and station back on air.

SDR/Metppf retrieved and merged in later in day.

Kh2oV at ~0v however we had lots of rain overnight and drizzle
	while there so we didn't consider it a problem.

DCOM332 - EVE problem?
	after 'starting' when replacing goesid (ie from s05_0110.xmt
	to s05_0120.xmt) the bootup failed claiming:
	'unable to open krypton.dat'
	even though it was there.   Played around a bit and had
	to perform a 'reboot' using the hard-pulled reset line low.
	That worked.
	ie teardown note: this cpu may have some intermittent on
	the ramdisk.  I'll slap in a note on the notes.
121: Site_visit, Site 5, Wed 15-Oct-2003 09:27:23 EDT, St 5: routine
10/14 13:35-15:00

Krypton:Lowered mast, cleaned Kr, ~400mv before; ~800mv after

Rad:	Cleaned Soldn and Netrad

Tsfc:	cleaned out a bit of spider web inside viewing port

Soil:	Trime measurement made, prongs would only penetrate 3/4 way
	Trime = 13.2
	Corer would only penetrate 4 cm. Core taken 2cm - 4 cm

Alignments:
Boom angles sighted with Datascope but as Battery low sighting compass used
 RMY boom = 187 degrees, Sonic boom = 97 (looking at it) degrees

Data:	Metppf and SDR data downloaded

RG:	Manual test OK, gauge was clean

GOES:	time OK.
	Did not replace GOES transmitter because backplane mounting plate
	doesn't have nutcerts for SE120 and it has been working OK with
	the old SE110 unit anyway.

117: Site_visit, Site 5, Mon 13-Oct-2003 17:13:02 EDT, St5: kh2oV checked
Arrive 14:35
Purpose: check for low kh2oV which had been seen in plots.

Did not have the spare GOES transmitter and merely checked Kr voltage = 500mv.
Decided against lowering mast to clean Kr. Will return soon



105: Site_visit, Site 5, Tue 07-Oct-2003 17:06:11 EDT, Site 5 visit, Oct 7
Site 5 visit, Oct 7, SPO & KDH
Arrived 1456 EDT

1502 mast down
cleaned Krypton 780-1050mV
1506 mast up

1520 TRIME 21.7, 17.4, 15.6%
soil core taken

cleaned rads
copy metppf
grabbed SDR

reboot






























81: Site_visit, Site 5, Mon 29-Sep-2003 18:10:52 EDT, Site 5 visit, Sept 29
Site 5 visit, Sept 29, SPO & KDH
Arrived 1454 EDT; brkn Cu

Cleaned krypton 380mV -> 700
Mast down 1504->1508

Grabbed SDR.
Copied metppf

1515: Trime: 30.0; 18.5; 29.7%
1515: took core

Cleaned rad

Rain gauge calibration 1458->1558 (255tips -> 358 = 103 tips)

GARMIN: 42deg 10.584'N; 73deg 49.577'W; 108m (acc 20ft)

1558 left site




67: Site_visit, Site 5, Thu 25-Sep-2003 16:47:34 EDT, Site 5 visit, Sept 24
Site 5 visit, Sept 24, SPO and Tom
Arrived 1405 EDT

kh2o  s/n 1394, output = 661  mV
copied kry1394s.dat to EVE and /card
downloaded metppf files and swapped fastout media

soil sample collected @ 1428 EDT, 3-6 cm
Qsoil = 30.4%
Trime = 11.4%	used pilot tool
  	20.3%	used pilot tool
	24.5%	no pilot tool

cleaned radiometers at 1448

SPO: GOES swap
~1420 "Bird' reports 7.3 W transmit
changed microstrip antenna; now 9.3 w
spoke w/ John M.; decided to swap SE110 1005 with 1006
1442: now 10.2 W forward, 0.2 W back
seems better; time will tell

mast down 1500 EDT, up 1506
cleaned kh2o, output = 665 mV


54: Site_visit, Site 5, Mon 22-Sep-2003 17:26:29 EDT, Site 5 visit, Sept 20
Visit to Site 5, Sept 20, Kurt
Arrived 1716 GMT
kh2o = 202 mV; all data look okay

Having trouble copying latest metppf file, metppf.263 to /card;
EVE continues to reboot (which automatically copies latest metppf
file anyway).  Did a dir /card and EVE again rebooted;
problem with PCMCIA card?
Tried a second time for dir /card; another reboot.
Downloaded local storage and now EVE is behaving normally again.
I can do a dir /card and there is no reboot.
Grabbed fastout.

Qsoil = 27.4% at 1746 GMT
Soil sample, 3-6 cm depth at 1749 GMT
Trime samples at 1751:	18.1%	used pilot tool
			22.1%	used pilot tool
			20.6%	used pilot tool

Cleaned radiometers at 1800 GMT


47: Site_visit, Site 5, Thu 18-Sep-2003 08:59:20 EDT, Site 5 visit; Sept 17
Visit to site 5, Sept 17; Kurt and Tom
Arrived 1245 EDT
Lowered mast 1259 EDT; Raised 1320 EDT

KH2O output = 518 mV; cleaned; output = 530 mV

Baro s/n = S0610001
CSAT s/n = 0537
RMY BoomDirection = 187 deg

Downloaded fastout and local data storage
Cleaned radiometers

Collected soil sample (3-6 cm) @ 1327 EDT
Trime =	23.2 %vol	using pilot tool
	14.4 %vol	w/out pilot tool
	20.8 %vol	using pilot tool
	25.7 %vol	w/out pilot tool


31: Site_visit, Site 5, Fri 12-Sep-2003 16:22:25 EDT, Visit: kh2o cleaned...
9/11/03 20:30Z
Kh2o 1394 cleaned:
	was 'pegged' at about 16.6mV
	cleaned with dry qtip and it came back up to ~250mV
metppf / SDR data uploaded

29: Site_visit, Site 5, Fri 12-Sep-2003 16:09:02 EDT, Site Visit: CARD problem, Kh20 installed, HFT fixed
9/07/03 ~23:45Z
	Kh20 2394 Installed
	PCMCIA /CARD: causing some difficulty at this site during boot.
		May be a problem with the board or the media.
	HFT wiring reversed due to sensor installed in ground 'upside-down'
19: Loggers, Site 5, Mon 11-Aug-2003 14:53:57 MDT, Site 5 CR10X calibration file
Q99259-A+       NET     00      9.07
A-      x       01      11.17
x       PygIn   02      0.0
x       PygOut  03      0.0
x       PspIn   04      0.0
x       Psput   05      0.0
H983211 HFT     06      36.4
PY10836 SlDn    07      -93.2
Tcin-A0 x       10      0.0
A1      x       11      0.0
A2      x       12      0.0
A3      x       13      0.0
A4      x       14      0.0
Tdin-A0 x       15      0.0
A1      x       16      0.0
A2      x       17      0.0
A3      x       18      0.0
A4      x       19      0.0
Tcot-A0 x       20      0.0
A1      x       21      0.0
A2      x       22      0.0
A3      x       23      0.0
A4      x       24      0.0
Tdot-A0 x       25      0.0
A1      x       26      0.0
A2      x       27      0.0
A3      x       28      0.0
A4      x       29      0.0
STP98032-A0     Tsoil   30      -241.435698
A1      x       31      232.471439
A2      x       32      9.245877
x-A0    LevelX  33      0.0
A1      x       34      0.0
A2      x       35      0.0
A3      x       36      0.0
A4      x       37      0.0
A5      x       38      0.0
x-A0    LevelY  39      0.0
A1      x       40      0.0
A2      x       41      0.0
A3      x       42      0.0
A4      x       43      0.0
A5      x       44      0.0
011-A0  CS615   45      -.187
A1      x       46      .037
A2      x       47      .335
TP01    timeout 50      180
levels  x       98      0
baud    x       99      1